OpenStudy (anonymous):

Chemistry Review Help! If 1.00 kilograms of coal is burned, approximately 3.00 x 107 joules (J) of energy is released. How many tons of coal, when burned, would produce approximately 9.00 x 1016 J? 1 ton = 907 kg 3.0 x 109 tons 3.3 x 106 tons 3.7 x 10-13 tons 3.3 x 10-10 tons

OpenStudy (anonymous):

@aaronq @dan815

OpenStudy (anonymous):

@surjithayer @KyanTheDoodle @Jamierox4ev3r @xapproachesinfinity @inkyvoyd @JoannaBlackwelder

OpenStudy (anonymous):

\[\frac{ 9.00 \times 10^{16} }{ 3.00 \times 10^{7} }=3 \times 10^{16-7} kg=3 \times 10^9 kg\] divide by 907 to convert in tons.

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

@surjithayer can you help me with another question? I just posted it a few minutes ago